A homogeneous triangular lamina has the vertices (0, 0), (3, 0) and (0, 3). Obtain the coordinates of the centre of mass of the lamina.


1
Expert's answer
2021-06-09T08:06:23-0400

"x_c=\frac{x_1+x_2+x_3}{3}=\frac{0+3+0}{3}=1"


"y_c=\frac{y_1+y_2+y_3}{3}=\frac{0+0+3}{3}=1" . So,


the coordinates of the centre of mass of the lamina "(1,1)" . Answer